7 6 5 CtN 5. On the given axes, show the following vector addition using arrows: where A = 3£+29 and B =-8+29 What is the magnitude, C? n the given axes, show the following result-
Answer #1

Vector C = (3-1) x^ + (2+2)y^ = 2x^ + 4y^

Now magnitude of C is |C| = √ (4+16 ) = √20 = 4.47